Use a suitable identity to get each of the following products.
ย ( ๐Ÿ/๐Ÿ ๐’™ + ๐Ÿ‘/๐Ÿ’ ๐’š)( ๐Ÿ/๐Ÿ ๐’™ + ๐Ÿ‘/๐Ÿ’ ๐’š)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

( ๐Ÿ/๐Ÿ ๐’™ + ๐Ÿ‘/๐Ÿ’ ๐’š)( ๐Ÿ/๐Ÿ ๐’™ + ๐Ÿ‘/๐Ÿ’ ๐’š)= (๐Ÿ/๐Ÿ๐’™+๐Ÿ‘/๐Ÿ’๐’š)^๐Ÿ
= (x^2/4) + (9y^2/16) + (3xy/4)
Using (a+b)^2= a^2+ b^2+ 2ab

Thanks